[figure: 9.7 cm]

Magnetic properties of RCuIn (R = Gd − Ho) compounds

M. Bałanda^{a}, V. Ivanov^{b} and A. Szytuła^{c}

aThe H. Niewodniczański Institute of Nuclear Physics, Radzikowskiego 152, 31-342

Kraków, Poland

bGeneral Physics Institute, Academy of Science, Vavilov Street 38, 117942 Moscow,

Russia

cInstitute of Physics, Jagiellonian University, Reymonta 4, 30-059 Kraków, Poland

AC susceptibility and DC magnetization measurements were performed for the RCuIn (R = Gd−Ho) series both in the paramagnetic and the ordred state. All compounds are antiferromagnets with the Néel temperatures equal 24 K for R = Gd, 28 K for R = T b, 6 K for R = Dy and 5.4 K for R = Ho. At 4.2 K the magnetization curves for R = Gd and T b are linear function of the external magnetic field while for R = Dy and Ho have the paramagnetic character. Above the Néel temperature the reciprocal magnetic susceptibilities obey the Curie - Weiss law with the paramagnetic Curie temperatures equal +4 K for R = Gd, -27 K R = T b, -10 K for R = Dy and -8 K for R = Ho and the effective magnetic moments near to the free R^{3+} ion values. The obtained results are compared with the data for the isostructural RT In (R = P d, Au) compounds.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : A. Szytuła

Address for correspondence : Institute of Physics, Jagiellonian University, Reymonta 4, 30-059 Kraków, Poland

Email address : szytula@if.uj.edu.pl